Mr. Kamal is an owner of the communication center at Y junction and uses four word processors and two typewriters to produce reports. The marginal product of a typewriter is 50 pages per day, and the marginal product of a word processor is 500 pages per day. The rental price of a typewriter is Rs.1 per day, whereas the rental price of a word processor is Rs.50 per day. Is Mr. Kamal utilizing type writers and word processors in a costminimizing manner? Which service process is more profitable? Type writer or word processing? Justify your answer


1
Expert's answer
2021-08-13T13:57:25-0400

Solution:

The cost-minimizing condition occurs when the marginal product of labor divided by the wage rate (w) is equal to the marginal product of capital divided by the rental price of capital (r).

"\frac{MP_{L} }{w} = \frac{MP_{K} }{r}"

Let MPL be the marginal product of typewriter.

Let MPK be the marginal product of word processors.

MPL = 50

W = 1


"\frac{MP_{L} }{w} =\frac{50}{1} = 50"

MPK = 500

R = 50


"\frac{MP_{K} }{r} =\frac{500}{50} = 10"

50 > 10


"\frac{MP_{L} }{w}" > "\frac{MP_{K} }{r}"

 

Mr. Kamal is not utilizing typewriters and word processors in a cost-minimizing manner. This is because the marginal product of typewriters divided by wage rate and the marginal product of word processors divided by the rental price is not equal.

For cost-minimization to be attained, they both must be equal.

 

The service process that is more profitable is a typewriter. This is because when the marginal product is higher or increasing, the total product increases at an increasing rate, and thus more profits are realized.